### v1

@@ Abstract
Foundation models trained with self-supervised learning (SSL) on large-scale histological images have significantly accelerated the development of computational pathology. These models can serve as backbones for region-of-interest (ROI) image analysis or patch-level feature extractors in whole-slide images (WSIs) based on multiple instance learning (MIL). Existing pathology foundation models (PFMs) are typically pre-trained on Hematoxylin-Eosin (H&E) stained pathology images. However, images with special stains, such as immunohistochemistry, are also frequently used in clinical practice. PFMs pre-trained mainly on H&E-stained images may be limited in clinical applications involving special stains. To address this issue, we propose StainNet, a specialized foundation model for special stains based on the vision transformer (ViT) architecture. StainNet adopts a self-distillation SSL approach and is trained on over 1.4 million patch images cropping from 20,231 publicly available special staining WSIs in the HISTAI database. To evaluate StainNet, we conduct experiments on an in-house slide-level liver malignancy classification task and two public ROI-level datasets to demonstrate its strong ability. We also perform few-ratio learning and retrieval evaluations, and compare StainNet with recently larger PFMs to further highlight its strengths. We have released the StainNet model weights at: https://huggingface.co/JWonderLand/StainNet.
@@
Special Staining, foundation model, computational pathology.1 Introduction
@@
Most existing PFMs are pre-trained on H&E-stained histological images, which are routinely used and easily accessible in clinical workflows. As a result, these PFMs typically perform well with H&E-stained images. For example, by directly analyzing expert-annotated H&E-stained ROIs, PFMs can be used to identify tumor malignancy or microenvironment patterns [lin2025tumor]. Additionally, when H&E-stained resection or biopsy WSI are cropped into patches, PFMs can extract these into high-dimensional features, which are used in weakly-supervised learning tasks based on MIL [xu2025multiple], such as cancer screening or biomarker detection [campanella2025real]. However, there are other staining techniques in clinical practice. For example, by using antibodies that bind to proteins in tissue cells, immunohistochemistry (IHC) slides are widely used to visualize the presence, location, and expression levels of specific proteins, which are commonly utilized for precision cancer diagnosis, disease subtyping, and drug selection. Due to significant differences in nuclear color, tissue patterns, and texture compared to H&E staining, directly transferring existing PFMs to special-stained images for representation may be limited in clinical performance [li2024unlocking].
To address this issue, we propose StainNet, a ViT-based PFM specifically pre-trained on special-staining images. Specifically, to train StainNet, we first select 20,231 non-H&E-stained WSIs from the publicly available large-scale HISTAI database [nechaev2025histai]. For each WSI, we randomly crop up to 100 tissue-containing patch images to create our pre-training dataset. Next, to develop StainNet, we utilize DINO [caron2021emerging], a lightweight self-supervised learning framework specifically designed for ViT models, enabling StainNet to learn task-agnostic morphological information from a large collection of unlabeled special staining images. Finally, to evaluate StainNet, we construct downstream task datasets at both the ROI and WSI levels and compare its performance with existing pathology foundation models. Through this work, we aim to explore the adaptability of PFMs in clinical tasks involving special-stained images and call for the broader exploration of computational pathology in more diverse clinical pathology image datasets.
@@ 2.1 Self-supervised PFMs
Foundation models are accelerating the advancement of computational pathology. By performing SSL on large-scale unlabeled pathology images, models can obtain histology-specific weights, which significantly improve their performance on downstream pathology tasks. For instance, CTransPath [wang2022transformer] uses a MoCov3-based SSL approach [chen2021empirical] and is trained on 15.6 million image patches from 25 anatomical sites to demonstrate its ability in H&E staining image retrieval. HIPT [chen2022scaling] and Lunit [kang2023benchmarking] adopt DINO [caron2021emerging], validating the feasibility of pre-training ViT models on histological images. Given the easy scalability of ViT parameters, recent works have aimed to train on larger datasets with larger ViT models. Examples include UNI [chen2024towards], Virchow [vorontsov2024foundation], and Prov-Gigapath [xu2024whole], which aim to diversify the pre-training dataset with more varied pathology images and leverage more computational resources to build stronger general pathology representation models. Additionally, some studies explore SSL in conjunction with special staining, such as PathoDuet [hua2024pathoduet], which utilizes registered H&E and IHC images to help the model learn cross-staining features, aiding the transfer to more advanced nuclear information.
2.2 Special staining in clinical pathology
@@
H&E staining is the most commonly used histological staining method, providing basic structural information by staining the cell nuclei and cytoplasm. However, H&E staining does not fully reveal certain cell features or tissue details, especially in the detection of immune responses or protein expression. As a result, several special staining methods have been developed to address these limitations [gridley1957manual]. Common special staining methods, such as IHC, Masson’s trichrome, and acid phosphatase staining, can highlight specific proteins, cell subtypes, or extracellular matrix components. For example, IHC utilizes antibodies to bind to target proteins, making it a valuable tool for cancer diagnosis and disease subtyping [magaki2018introduction]; Masson’s trichrome staining is employed to visualize elastic fibers in tissues, particularly in fibrosis-related research [lefkowitch2006special]. Recently, the development of computational pathology has accelerated research into the application of computer vision models to these special staining images, such as lymphocyte detection in IHC images [swiderska2019learning] and the use of multiple special stains for kidney biopsy evaluation [jayapandian2021development]. Some studies have also explored the direct use of generative models to convert H&E images into special stains to improve diagnostic accuracy [de2021deep, bai2023deep, pati2024accelerating, yan2023unpaired]. However, these special staining images differ significantly from H&E staining in terms of staining mechanisms, texture features, and visual appearance, making it difficult for large-scale pre-trained models based on general H&E staining images to capture the specific features of special staining images [li2024unlocking].
@@ 3 Methodology
We develop and evaluate our StainNet through data collection, model pre-training, and downstream testing. The entire process is illustrated in Figure 1.
@@ 3.1 Data collection
We use the publicly available large-scale WSI database HISTAI [nechaev2025histai] as our pre-training data. The original HISTAI dataset contains 112,801 WSIs, of which 20,265 are special stains. We first apply the OTSU thresholding algorithm [otsu1975threshold] to exclude 34 special staining WSIs that can not extract foreground tissue, resulting in 20,231 WSIs. Then, for each WSI, we perform non-overlapping pixel sliding window operations on the foreground tissue, and randomly crop 100 image patches (if a WSI contains fewer than 100 patches, all the patches will be selected). In total, we obtain 1,418,938 patches for subsequent StainNet pre-training.
@@ 3.2 Model pre-training
Our proposed StainNet adopts the SSL method DINO [caron2021emerging] as the pre-training strategy to let the ViT model learn robust feature representations from special-staining pathology images without manual annotations. DINO consists of two networks with the same architecture: a student network optimized via gradient descent and a teacher network updated using an exponential moving average (EMA) of the student weights. We follow the original DINO approach for data augmentation and multi-crop strategies. Specifically, we generate a set of global and local views for each pathology image, and the student takes all views as input, while the teacher only receives the global views. By minimizing the cross-entropy loss between the output probability distributions of the two networks, the student is encouraged to infer global tissue context solely from local information. In addition, centering and sharpening operations are applied to the teacher outputs to prevent mode collapse. Figure 2 illustrates some examples of original images and their augmented versions. The detailed hyperparameters are provided in the implementation.
@@ 4.1 Downstream datasets and implementation
We evaluate StainNet on one private WSI task NTUH-Ki67-Liver, and two publicly available ROI tasks BCI [liu2022bci] and MIST [li2023adaptive]. The NTUH-Ki67-Liver 3-class dataset consists of 46 Ki67 benign liver slides, 43 Ki67 hepatocellular carcinoma slides, and 50 Ki67 normal liver slides. All slides are collected from the Affiliated Hospital of Nantong University and scanned using the SQS-120P scanning system from Shenzhen Shengqiang Technology Co., Ltd. The annotations are performed and verified at the slide level by two pathologists. The BCI dataset is a collection of 4,870 paired H&E-IHC images showing HER2 biomarker expression. We use the IHC images and categorize them into four HER2 expression levels: 0, 1+, 2+, and 3+, with 240, 1,153, 2,142, and 1,355 samples, respectively. The MIST dataset is a large-scale H&E-IHC paired image dataset. We use the IHC images corresponding to four biomarkers: 5,642 HER2, 5,361 Ki67, 5,153 ER, and 5,139 PR to construct a 4-class dataset for evaluating the ability of PFMs to identify biomarkers across different staining conditions.
We use the AdamW optimizer with a learning rate of to pre-train StainNet for 100 epochs with a batch size of 256. For multi-cropping, global crops are sampled using a random resize scale in the range of 0.4-1.0, and 8 local crops are generated with a scale in the range of 0.05-0.4. For EMA design, we increase teacher momentum linearly from 0.9995 to 1.0 throughout training. The weight decay is scheduled from 0.04 to 0.4, and the teacher temperature is set to 0.04 without warm-up and last layer freezing. We use ViT-Small, pre-trained on ImageNet datasets [deng2009imagenet], as the initial backbone and weights for our StainNet. We also enable normalization of the last layer to stabilize the training process. The experiment is conducted on a single NVIDIA RTX PRO 6000 96GB GPU with mixed-precision and takes approximately 3.04 days to complete. We select the final epoch checkpoint as the StainNet weights used in all subsequent experiments.
For downstream fine-tuning, we use AdamW with a learning rate of and a weight decay of . We set 20 epochs with a batch size of 1 for the WSI task and 20 epochs with a batch size of 128 for the ROI task. Early stopping is set to 5 epochs for both functions. For classification tasks, we save the fine-tuned weights and report accuracy, balanced accuracy, AUC, and F1-score from each epoch that achieves the best balanced accuracy. For retrieval tasks, we use cosine similarity for image matching and report recall@K and mAP@K, where K ranges from 1 to 20. All experiments are conducted using stratified 5-fold cross-validation and full-precision computing on a single NVIDIA RTX 3090 24GB GPU. All baseline methods are trained using the same settings.
@@
| Bal acc. | ABMIL | SiMLP | TransMIL | WiKG | AMDMIL | S4MIL | Overall |
| ResNet-50 | 0.7980 | ||||||
| CTransPath | 0.7940 | ||||||
| PathoDuet | 0.6957 | ||||||
| HIPT | 0.7943 | ||||||
| Lunit | 0.8470 | ||||||
| StainNet (Ours) | 0.8769 | ||||||
| Patch encoder | MIL encoder | ||||||
| AUC | ABMIL | SiMLP | TransMIL | WiKG | AMDMIL | S4MIL | Overall |
| ResNet-50 | 0.9019 | ||||||
| CTransPath | 0.9127 | ||||||
| PathoDuet | 0.8382 | ||||||
| HIPT | 0.8964 | ||||||
| Lunit | 0.9275 | ||||||
| StainNet (Ours) | 0.9492 | ||||||
| Patch encoder | MIL encoder | ||||||
| F1-score | ABMIL | SiMLP | TransMIL | WiKG | AMDMIL | S4MIL | Overall |
| ResNet-50 | 0.7931 | ||||||
| CTransPath | 0.7671 | ||||||
| PathoDuet | 0.6668 | ||||||
| HIPT | 0.7801 | ||||||
| Lunit | 0.8409 | ||||||
| StainNet (Ours) | 0.8706 | ||||||
4.2 Comparison with PFMs of similar parameter scale

### v2

@@ Abstract
Foundation models trained with self-supervised learning (SSL) on large-scale histological images have significantly accelerated the development of computational pathology. These models can serve as backbones for region-of-interest (ROI) image analysis or patch-level feature extractors in whole-slide images (WSIs) based on multiple instance learning (MIL). Existing pathology foundation models (PFMs) are typically pre-trained on Hematoxylin-Eosin (H&E) stained pathology images. However, images such as immunohistochemistry (IHC) and special stains are also frequently used in clinical practice. PFMs pre-trained mainly on H&E-stained images may be limited in clinical applications involving these non-H&E images. To address this issue, we propose StainNet, a collection of self-supervised foundation models specifically trained for IHC and special stains in pathology images based on the vision transformer (ViT) architecture. StainNet contains a ViT-Small and a ViT-Base model, both of which are trained using a self-distillation SSL approach on over 1.4 million patch images extracted from 20,231 publicly available IHC and special staining WSIs in the HISTAI database. To evaluate StainNet models, we conduct experiments on three in-house slide-level IHC classification tasks, three in-house ROI-level special stain and two public ROI-level IHC classification tasks to demonstrate their strong ability. We also perform ablation studies such as few-ratio learning and retrieval evaluations, and compare StainNet models with recent larger PFMs to further highlight their strengths. The StainNet model weights are available at https://github.com/WonderLandxD/StainNet.
@@
Non-H&E images, foundation model, computational pathology.1 Introduction
@@
Most existing PFMs are pre-trained on H&E-stained histological images, which are routinely used and easily accessible in clinical workflows. As a result, these PFMs typically perform well with H&E-stained images. For example, by directly analyzing expert-annotated H&E-stained ROIs, PFMs can be used to identify tumor malignancy or microenvironment patterns [lin2025tumor]. Additionally, when H&E resection or biopsy WSIs are cropped into patches, PFMs can extract these into high-dimensional features, which are used in weakly-supervised learning tasks based on MIL [xu2025multiple], such as cancer screening or biomarker detection [campanella2025real]. However, there are other staining in clinical practice. For example, by using antibodies that bind to proteins in tissue cells, IHC slides are widely used to visualize the presence, location, and expression levels of specific proteins, which are commonly utilized for precision cancer diagnosis, disease subtyping, and drug selection. Due to significant differences in nuclear color, tissue patterns, and texture compared to H&E staining, directly transferring existing PFMs to these non-H&E images for representation may be limited in clinical performance [li2024unlocking].
To address this issue, we propose StainNet, a collection of PFMs specifically pre-trained on IHC and special stain images. Specifically, to train StainNet, we first select 20,231 non-H&E WSIs from the publicly available large-scale HISTAI database [nechaev2025histai]. For each WSI, we randomly crop up to 100 tissue-containing patch images to create our pre-training dataset. Next, to develop StainNet, we utilize DINO [caron2021emerging], a lightweight self-supervised learning framework specifically designed for ViT models, enabling StainNet to learn task-agnostic morphological information from a large collection of unlabeled special staining images. Finally, to evaluate StainNet, we construct downstream task datasets at both the ROI and WSI levels and compare textcolorredthe performance with existing pathology foundation models. Through this work, we aim to explore the adaptability of PFMs in clinical tasks involving non-H&E images and call for the broader exploration of computational pathology in more diverse clinical pathology image datasets.
@@ 2.1 Self-supervised PFMs
Foundation models are accelerating the advancement of computational pathology. By performing SSL on large-scale unlabeled pathology images, models can obtain histology-specific weights, which significantly improve their performance on downstream tasks. For instance, CTransPath [wang2022transformer] uses a MoCov3-based SSL approach [chen2021empirical] and is trained on 15.6 million image patches from 25 anatomical sites to demonstrate its ability in H&E staining image retrieval. HIPT [chen2022scaling] and Lunit [kang2023benchmarking] adopt DINO [caron2021emerging], validating the feasibility of pre-training ViT models on histological images. Given the easy scalability of ViT parameters, recent works have aimed to train on larger datasets with larger ViT models. Examples include UNI [chen2024towards], Virchow [vorontsov2024foundation], and Prov-Gigapath [xu2024whole], which aim to diversify the pre-training dataset with more varied pathology images and leverage more computational resources to build stronger general pathology representation models. Additionally, some studies explore SSL in conjunction with special staining, such as PathoDuet [hua2024pathoduet], which utilizes registered H&E and IHC images to help the model learn cross-staining features, aiding the transfer to more advanced nuclear information.
2.2 IHC and Special staining in clinical pathology
@@
H&E staining is the most commonly used histological staining method, providing basic structural information by staining the cell nuclei and cytoplasm. However, H&E staining does not fully reveal certain cell features or tissue details, especially in the detection of immune responses or protein expression. To address these limitations, several staining methods such as IHC, Masson’s trichrome, and acid phosphatase staining are developed to highlight the specific proteins, cell subtypes, and extracellular matrix components [gridley1957manual]. For example, IHC utilizes antibodies to bind to target proteins, making it a valuable tool for cancer diagnosis and disease subtyping [magaki2018introduction]; Masson’s trichrome staining is employed to visualize elastic fibers in tissues, particularly in fibrosis-related research [lefkowitch2006special]. Recently, the development of computational pathology has accelerated research into the application of computer vision models to these special staining images, such as lymphocyte detection in IHC images [swiderska2019learning] and the use of multiple special stains for kidney biopsy evaluation [jayapandian2021development]. Some studies have also explored the direct use of generative models to convert H&E images into these stains to improve diagnostic accuracy [de2021deep, bai2023deep, pati2024accelerating, yan2023unpaired]. However, these non-H&E images differ significantly from H&E staining in terms of staining mechanisms, texture features, and visual appearance, making it difficult for large-scale pre-trained models based on general H&E staining images to capture the specific features of special staining images [li2024unlocking].
@@ 3 Methodology
We develop and evaluate our StainNet models through data collection, model pre-training, and downstream testing. The entire process is illustrated in Figure 1.
@@ 3.1 Data collection
We use the publicly available large-scale WSI database HISTAI [nechaev2025histai] as our pre-training data. The original HISTAI dataset contains 112,801 WSIs, of which 20,265 are non-H&E stains. We first apply the OTSU thresholding algorithm [otsu1975threshold] to exclude 34 special staining WSIs that can not extract foreground tissue, resulting in 20,231 WSIs. Then, for each WSI, we perform non-overlapping pixel sliding window operations on the foreground tissue, and randomly crop 100 image patches (if a WSI contains fewer than 100 patches, all the patches will be selected). In total, we obtain 1,418,938 patches for subsequent StainNet pre-training. More pre-training data details are provided in Appendix A.
@@ 3.2 Model pre-training
Our proposed StainNet models adopt the SSL method DINO [caron2021emerging] as the pre-training strategy to let the ViT model learn robust feature representations from IHC and special stain pathology images without manual annotations. DINO consists of two networks with the same architecture: a student network optimized via gradient descent and a teacher network updated using an exponential moving average (EMA) of the student weights. We follow the original DINO approach for data augmentation and multi-crop strategies , which are provided in Appendix B. Specifically, we generate a set of global and local views for each pathology image, and the student takes all views as input, while the teacher only receives the global views. By minimizing the cross-entropy loss between the output probability distributions of the two networks, the student is encouraged to infer global tissue context solely from local information. In addition, centering and sharpening operations are applied to the teacher outputs to prevent mode collapse. The detailed hyperparameters are provided in the implementation and Appendix B.
@@ 4.1 Downstream datasets and implementation
We evaluate StainNet on three private IHC WSI tasks, NTUH-Ki67-Liver, P53-UCEC, and MLH1-UCEC, two public available IHC ROI tasks, BCI [liu2022bci] and MIST [li2023adaptive], and three in-house special stain ROI tasks, Glomerulus-Masson, Glomerulus-PAS, and Glomerulus-PASM, which are derived as subsets from our previous work [he2025unveiling]. The NTUH-Ki67-Liver 3-class dataset consists of 46 Ki67 benign liver slides, 43 Ki67 hepatocellular carcinoma slides, and 50 Ki67 normal liver slides. The P53-UCEC 5-class dataset consists of 139 P53 wild-type, 38 P53 mutant, 10 P53 null, 14 P53 subclonal, and 17 P53 endometrial carcinoma slides. The MLH1-UCEC 3-class dataset consists of 30 MLH1 deficient, 155 MLH1 proficient, and 10 slides with partial loss of MLH1 expression from endometrial carcinoma. All slides are collected from the Affiliated Hospital of Nantong University and scanned using the SQS-120P scanning system from Shenzhen Shengqiang Technology Co., Ltd. The annotations are performed and verified at the slide level by two pathologists. The BCI dataset is a collection of 4,870 paired H&E-IHC images showing HER2 biomarker expression. We use the IHC images and categorize them into four HER2 expression levels: 0, 1+, 2+, and 3+, with 240, 1,153, 2,142, and 1,335 samples, respectively. The MIST dataset is a large-scale H&E-IHC paired image dataset. We use the IHC images corresponding to four biomarkers: 5,642 HER2, 5,361 Ki67, 5,153 ER, and 5,139 PR to construct a 4-class dataset for evaluating the ability of PFMs to identify biomarkers across different staining conditions. The Glomerulus-Masson dataset consists of 200 normal, 57 slight, 112 moderate, and 113 severe Masson-stained glomeruli images. The Glomerulus-PAS dataset consists of 1200 normal, 1129 slight, 479 moderate, and 379 severe PAS-stained glomeruli images. The Glomerulus-PASM dataset consists of 200 normal, 76 slight, 135 moderate, and 87 severe PASM-stained glomeruli images.
We further evaluate StainNet models on two public H&E-stained ROI tasks, CRC-100K [kather2019predicting] and KatherMS [kather2019deep], as well as one WSI task, PANDA-Karo. [bulten2022artificial], to investigate its capability on H&E image analysis. CRC-100K consists of nine categories: adipose, background, debris, lymphocytes, mucus, smooth muscle, normal colon mucosa, cancer-associated stroma, and colorectal adenocarcinoma epithelium. We perform label-stratified splitting on the official NCT-CRC-HE-100K dataset with a ratio of 0.8:0.2 to construct the train–val set, and use CRC-VAL-HE-7K as the test set. KatherMS contains two categories: microsatellite stable and microsatellite instable. We similarly conduct label-stratified splitting on the official training set with a ratio of 0.8:0.2 to obtain the train–val set, and use the official test set for evaluation. PANDA-Karo. consists of 5455 biopsy WSIs from Karolinska Institute, including 1924 G0, 1814 G1, 668 G2, 317 G3, 481 G4, and 251 G5. We label-stratify the train-val-test fold with ISUP grading ground truth of 0.5:0.2:0.3 for experiments.
@@
| NTUH-Ki67-Liver | ABMIL | SiMLP | TransMIL | WiKG | AMDMIL | S4MIL | Overall |
| ResNet-50 [he2016deep] | 0.7980 | ||||||
| CTransPath [wang2022transformer] | 0.7940 | ||||||
| PathoDuet [hua2024pathoduet] | 0.6957 | ||||||
| HIPT [chen2022scaling] | 0.7943 | ||||||
| Lunit [kang2023benchmarking] | 0.8470 | ||||||
| UNI [chen2024towards] | 0.8755 | ||||||
| PathOrchestra [yan2025pathorchestra] | 0.8367 | ||||||
| GPFM [ma2025generalizable] | 0.8630 | ||||||
| UNI-2 [chen2024towards] | 0.8833 | ||||||
| Virchow-2 [zimmermann2024virchow2] | 0.8895 | ||||||
| Prov-GigaPath [xu2024whole] | 0.8847 | ||||||
| H-optimus-0 [hoptimus0] | 0.8592 | ||||||
| StainNet-Small (Ours) | 0.8768 | ||||||
| StainNet-Base (Ours) | 0.8930 | ||||||
For DINO training, we use the AdamW optimizer with a learning rate of to pre-train StainNet for 100 epochs with a batch size of 256. For multi-cropping, global crops are sampled using a random resize scale in the range of 0.4-1.0, and 8 local crops are generated with a scale in the range of 0.05-0.4. For EMA design, we increase teacher momentum linearly from 0.9995 to 1.0 throughout training. The weight decay is scheduled from 0.04 to 0.4, and the teacher temperature is set to 0.04 without warm-up and last layer freezing. Our StainNet model family includes StainNet-Small and StainNet-Base, which are initialized with ImageNet-pre-trained [deng2009imagenet] ViT-Small and ViT-Base backbones, and are then further pre-trained using DINO. We also enable normalization of the last layer to stabilize the training process. The StainNet-Small experiment is conducted on a single NVIDIA RTX PRO 6000 96GB GPU with mixed-precision and takes approximately 3.04 days to complete. The StainNet-Base experiment is conducted on 8 NVIDIA A100 80GB GPU and tasks approximately 2.78 days to complete. For both StainNet-Small and StainNet-Base, we select the last epoch checkpoint as the final model weights.
For downstream fine-tuning, we use AdamW with a learning rate of and a weight decay of . We set 20 epochs with a batch size of 1 for the WSI task and 20 epochs with a batch size of 128 for the ROI task. For classification tasks, we save the fine-tuned model weights based on the best balanced accuracy, and also report accuracy, AUC, and F1-score from each fold. Early stopping is applied with a patience of 5 epochs, meaning that training is terminated if the balanced accuracy on the validation set does not improve for five consecutive epochs. For retrieval tasks, we perform image matching using cosine similarity and report Recall@K and mAP@K, where K ranges from 1 to 20. For NTUH-Ki67-Liver, P53-UCEC, MLH1-UCEC, BCI, and MIST, we adopt a label-stratified 5-fold train–val cross-validation setting for evaluation. For Glomerulus-Masson, Glomerulus-PAS, and Glomerulus-PASM, we follow our previous work [li2024unlocking] to employ a label-stratified train–val–test split of 0.5:0.2:0.3 with five random seeds. For H&E tasks, we use the train-val-test setting with three random seeds. We use a single NVIDIA RTX 3090 24GB GPU with full-precision computing to run all downstream tasks. For MIL and PFM baselines, we follow their official implementations, including hyperparameter settings and data preprocessing pipelines. Except for these model-specific configurations, all other downstream fine-tuning hyperparameters are kept identical to those used in StainNet.
4.2 Comparison results on WSI tasks
| Patch encoder | P53-UCEC | MLH1-UCEC | ||||||
| Acc. | Bal acc. | AUC | F1-score | Acc. | Bal acc. | AUC | F1-score | |
| ResNet-50 [he2016deep] | ||||||||
| CTransPath [wang2022transformer] | ||||||||
| PathoDuet [hua2024pathoduet] | ||||||||
| HIPT [chen2022scaling] | ||||||||
| Lunit [kang2023benchmarking] | ||||||||
| UNI [chen2024towards] | ||||||||
| PathOrchestra [yan2025pathorchestra] | ||||||||
| GPFM [ma2025generalizable] | ||||||||
| UNI-2 [chen2024towards] | ||||||||
| Virchow-2 [zimmermann2024virchow2] | ||||||||
| Prov-GigaPath [xu2024whole] | ||||||||
| H-optimus-0 [hoptimus0] | ||||||||
| StainNet-Small (Ours) | ||||||||
| StainNet-Base (Ours) | ||||||||